OpenStudy (anonymous):

Given m∠A = 9x + 1, m∠B=3x and m∠C = 11x - 5, find m∠A, m∠B and m∠C.

OpenStudy (anonymous):

is this in a triangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you know that the angles of triangle add up to 180degrees

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so, the sum of m∠A, m∠B and m∠C. = 180, because there is only \[x\] involved we can sum all the angle formulae that is A+B+C=180

OpenStudy (anonymous):

\[(9x+1) + 3x + (11x-5)=180\]

OpenStudy (anonymous):

combine like terms..solve for x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i'll be here if you want to check

OpenStudy (anonymous):

is the answer 8

OpenStudy (anonymous):

\[23x-4=180\] \[x=(180+4)/23\]\[x=184/23\] \[x=8\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so now, put x back in to each equation, what is angle A? B? C?

OpenStudy (anonymous):

A=73, B=24, C=83

OpenStudy (anonymous):

i figured it out thanks